Looking for the steps for the double exponential smoothing part of b. Please DO NOT use excel. Thank you. Q3. (Total 25 points) Goodyear Tire and Rubber Company is the ninth largest tire manufacturer in the world. Here are the sale
Answer #1

Double exponentially smoothed forecast (or trend-adjusted exponentially smoothed) forecast is calculated by the following formulas:

St+1 = \alpha*At + (1-\alpha)*St, where

Tt+1 = \beta*(St+1-St) + (1-\beta)*Tt

TAFt+1 = St+1+Tt+1

For year 2, S2 = .4*4887.9+(1-.4)*4867.9 = 4875.9

T2 = .2*(4875.9-4867.9)+(1-0.2)*201.5 = 162.8

TAF2 = 4875.9+162.8 = 5038.7

--------------

For year 3, S3 = .4*5065.4+(1-.4)*4875.9 = 4951.7

T3 = .2*(4951.7-4875.9)+(1-0.2)*162.8 = 145.5

TAF3 = 4951.7+145.5 = 5097.1

--------------

For year 4, S4 = .4*5535.6+(1-.4)*4951.7 = 5185.26

T4 = .2*(5185.26-4951.7)+(1-0.2)*145.5 = 163.03

TAF4 = 5185.26+163.03 = 5348.29

--------------

For year 5, S5 = .4*5730.8+(1-.4)*5185.26 = 5403.48

T5 = .2*(5403.48-5185.26)+(1-0.2)*163.03 = 174.07

TAF5 = 5403.48+174.07 = 5577.55

--------------

For year 6, S6 = .4*5499.7+(1-.4)*5403.48 = 5441.97

T6 = .2*(5441.97-5403.48)+(1-0.2)*174.07 = 146.95

TAF6 = 5441.97+146.95 = 5588.92

--------------

[Forecast for year 7 cannot be calculated using double-exponential smoothing method]